dm- 0 q n 0 hhd nco 1 tec e o 5 lgs Ce 3 e t rei eye h︶u es 5888 5169 0 5556 eo Cay 5
Answer #1

Find PV of withdrawal and deposits and equate them

PV deposit = A + A*(P/A, 10%, 7 yrs)

= A + A* 4.8684 = A *(1+ 4.8684) = A* 5.8684

PV withdrawl= (1000(P/A,10%,4)+200(P/G,10%,4))(P/F,10%,3)

=(1000X3.1698+200X4.3781)(0.75131)

=3039.43

Setting PV of withdrawl = PV of deposits

A x 5.8684 = 3039.43

A = 3039.43/5.8684

A = 518

Option(B)